Sol Summary – May 25th

Crew 281 Sol Summary Report 25 May 2023
Sol: 11
Summary Title: Geological miracles
Author’s name: Ritupriya Patil (XO) and Megan Kane (Commander)
Mission Status: Nominal
Sol Activity Summary:
This was another long EVA with a crew of three. The crew entered the airlock exiting at 0753 hrs. They first checked all the rovers and collected the hours and SoC information, and it was verified that all were securely plugged in. They then unplugged Opportunity and Curiosity and headed out for the first destination at 0758 for the intersection of Cow Dung and Watney Road. The area seemed like a former riverbed; the crew saw lots of sedimentary rocks possibly washed by the flowing water in the past. They also witnessed calcite compressions on the sides of the pathway. The crew walked no more than 600 m and explored. Ana (Crew Scientist) collected 7 rocks (1167 g) and 1 Tube of soil (72 g). Ritu (XO, EVA Lead) deployed the drone to capture aerial.
For lunch the crew had leftovers and then worked on mission and research summaries. Rachel accomplished her goal of 50 contacts on her ham radio set up across 2 nations – USA and Canada. KC did his routine checks along with inventory work. Megan prepped green hab for end of mission dismantling. Ana did the same for science dome and the RAM. Ritu worked on the daily reports and EVA planning for tomorrow, which will be the longest of all.
For dinner Megan prepared potato vegetable soup for dinner and then crew jumped on a call with the current Hi-SEAs crew called IKE-LOA. During the call, both crews exchanged their experience with research, facilities and showed around their respective habitats. It was a great way to connect with another crew doing similar mission, Hi-SEAs was on the Moon and MDRS crew was on Mars. This was also a great opportunity to network with other space explorers.
Look Ahead Plan: Tomorrow we will do 2-person, long range EVA to run some more drone payload tests. Next we will stop at Marble Ritual for images of scanlines that were measured on earlier EVAs. Last stop will be the Special Region in the north.
Anomalies in work: None
Weather: Winds
Crew Physical Status: Nominal
EVA: Completed 1 today
Reports to be filed:
Greenhab Report
EVA Report
EVA Request
Ops Report
Journalist Report
Photo of the Day

Support Requested:
Mission support is requested in case of storms if the antenna needs to come down or be retrieved.

Copyright © The Mars Society. All rights reserved. | Main Site